OpenStudy (anonymous):

Joe got a bonus of $1,600.20 based on a certain percent of his sales for the year, which was $22,860. What percent of his sales for the year was the bonus? A. 5.5% B. 7% C. 14% D. 15%

OpenStudy (ankit042):

@renae_ can you try solving this...where are you stuck?

OpenStudy (anonymous):

i dont know percentages

hero (hero):

Formula to use: \[\frac{\text{part}}{\text{whole}} = \frac{\%}{100}\] 1600.20 = part of 22,860, the whole and the percentage,x, is what we have to find so: \[\frac{1600.20}{22860} = \frac{x}{100}\] Cross multiply and solve.

hero (hero):

@renae_ Did you solve for x yet?

OpenStudy (anonymous):

its 7 (:

hero (hero):

Did you understand the formula?

OpenStudy (anonymous):

yes c:

hero (hero):

Great